# Your Image is Secretly the Last Frame of a Pseudo Video

### v2

@@ 2.3 Improving Image Generation via Pseudo Video Generation
With the concept of pseudo videos and their effectiveness in diffusion models, we are interested in the following open generic question in this paper: {adjustwidth}2em1.5em Is it possible to improve other types of image generative models by jointly modelling the distribution of the original image and its corresponding pseudo video which contains self-supervised information? The answer is affirmative. In this paper, we show empirical evidence of the advantages of pseudo videos on two types of generation models, namely improving VQVAE (Van Den Oord et al., 2017) (Section 3) and DDPM (Nichol & Dhariwal, 2021) (Section 4.2) with Phenaki (Villegas et al., 2022) and Video Diffusion (Harvey et al., 2022) trained on pseudo videos, respectively. Moreover, we provide theoretical arguments favouring the use of more expressive ways of creating pseudo videos in the autoregressive video generation framework (Section 4.1), beyond the first-order Markov strategy as typically used in the forward process of standard diffusion models.
@@ 3 Improved Reconstruction and Generation in VQ-VAE with Pseudo Videos
[…]
@@
Evaluation Metric. We compute Frechet Inception Distance (FID) (Heusel et al., 2017) with 50k samples to evaluate the quality of images, either from the last-frames of videos generated by video models trained on pseudo videos, or images generated by image models trained on the original target images.
@@
Results. Table 2 shows the last-frame FID of C-ViViT for reconstruction (we also report Peak Signal-to-noise Ratio (PSNR) for reconstruction in Appendix D) and that of VideoGPT and Phenaki for AR and masked generation on CIFAR10 and CelebA, respectively. The 1-frame results correspond to the performance of their image model counterparts (i.e., VQVAE for image reconstruction, and ImageGPT and MaskGit for AR and masked image generation, respectively). We observe that pseudo videos indeed help improve the training of the C-ViViT as the reconstruction quality of the last frame is significantly improved with a few more frames. We show reconstructed CIFAR10 and CelebA images from different C-ViViT models in Figures LABEL:fig:recon_cifar10 and LABEL:fig:recon_celeba, respectively. It can be seen that for both datasets, the reconstructed images trained with 1-frame models are over-smoothed and this issue is resolved by using pseudo videos which produce much sharper images. Quantitatively, reconstruction FID improves as more frames are used. Pseudo videos also improve the image generation performance compared to the 1-frame results. Interestingly, we see a diminishing return as we include more frames. For CelebA images, 18-frame pseudo videos help achieve the best image generation performance for both AR and latent masked generation, but 8-frame models achieve a comparable performance. For CIFAR10 images, 8-frame pseudo videos result in better image generation performance than 18-frame pseudo videos, which suggests the latent codes have a more complex prior distribution in order to reconstruct model pseudo videos with 18 frames well and therefore this prior is more difficult for VideoGPT or Phenaki to capture. In summary. while pseudo videos help improve the performance, the optimal number of frames may depend on the dataset and the augmentation strategy. This diminishing return is not a severe issue in practice since practitioners may prefer to improve the generation with just a few more pseudo frames to avoid introducing a high computational cost.
@@
Table 2: Last-frame FID of images produced by C-ViViT (reconstruction), VideoGPT (AR generation) and Phenaki (latent masked generation) trained on pseudo videos constructed from CIFAR10 and CelebA images. 1-frame results are obtained from their image counterparts VQVAE (reconstruction), ImageGPT (AR generation) and MaskGit (latent masked generation) trained on original CIFAR10 and CelebA images.
CIFAR10 CelebA
1-frame 8-frame 18-frame 1-frame 8-frame 18-frame
Reconstruction 84.25 13.81 11.26 24.62 5.72 2.27
AR Generation 91.65 54.60 69.23 32.98 30.19 28.08
Latent Masked Generation 89.78 35.50 47.65 27.34 16.87 16.66
@@ 4.1 Is First-order Markov Chain the Optimal Choice for Creating Pseudo Videos?
Moreover, the non-optimality of creating pseudo video via first-order Markov chain becomes clear: the first-order Markov data augmentation implies that $p(x_{T}|x_{T-1})=p(x_{T}|x_{T-1},x_{T-2})$ and consequently $\mathcal{L}_{2}^{*}=\mathcal{L}_{1}^{*}$. More specifically, for strict inequality in Eq 14, we need to avoid $p(x_{T}|x_{T-1})=p(x_{T}|x_{T-1},x_{T-2})$, which is equivalent to avoiding the use of either first-order Markov chain $x_{T}\rightarrow x_{T-1}\rightarrow x_{T-2}$ or $x_{T}\leftarrow x_{T-1}\rightarrow x_{T-2}$. This analysis is informative for us to design better pseudo videos, for example through data augmentation with higher-order Markov chains. We formalize the above informal reasoning into Theorem 4.1 and provide the formal proof in Appendix B.
@@
$\min_{\hat{x}_{T}^{(h)}}E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})}[||x_{T}-\hat{x}_{T}^{(h)}||_{2}^{2}]\leq\min_{\hat{x}_{T}^{(g)}}E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[||x_{T}-\hat{x}_{T}^{(g)}||_{2}^{2}],$ (15)
@@ 4.2 Experiments
[…]
@@ 6 Conclusion and Discussions
Appendix A Reproducibility Statement
To our knowledge, the official implementation of Phenaki (Villegas et al., 2022) has not been released upon submission of this paper. As a result, the experiments in Section 3 are based on the implementation from https://github.com/lucidrains/phenaki-pytorch. The experiments in Section 4.2 are based on the official implementation of Flexible Video Diffusion Model (FDM) (Harvey et al., 2022) and Improved DDPM (Nichol & Dhariwal, 2021), from https://github.com/plai-group/flexible-video-diffusion-modeling and https://github.com/openai/improved-diffusion.
Appendix B Proof of Theorem 4.1
@@ Appendix A Reproducibility Statement
To derive $\hat{x}_{T}^{(g)*}(x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})=\operatorname*{arg\,min}_{\hat{x}_{T}^{(g)}}E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[||x_{T}-\hat{x}_{T}^{(g)}||^{2}_{2}]$, we first compute the gradient,
$\begin{split}\nabla_{\hat{x}_{T}^{(g)}}E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[||x_{T}-\hat{x}_{T}^{(g)}||^{2}_{2}]&=E_{p(x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}\{\nabla_{\hat{x}_{T}^{(g)}}E_{p(x_{T}|x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[||x_{T}-\hat{x}_{T}^{(g)}||^{2}_{2}]\}\\ &=2E_{p(x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}\{E_{p(x_{T}|x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[x_{T}-\hat{x}_{T}^{(g)}]\}.\\ \end{split}$ (16)
@@
The minimum reconstruction error is obtained by plugging $\hat{x}_{T}^{(g)*}(x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})$ in $E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[||x_{T}-\hat{x}_{T}^{(g)}||^{2}_{2}]$,
$\begin{split}\min_{\hat{x}_{T}^{(g)}}E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[||x_{T}-\hat{x}_{T}^{(g)}||^{2}_{2}]&=E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[||x_{T}-E_{p(x_{T}|x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[x_{T}]||^{2}_{2}]\\ &=E_{p(x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[\text{Var}_{p(x_{T}|x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}(x_{T})].\end{split}$ (18)
@@
$\begin{split}\min_{\hat{x}_{T}^{(h)}}E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})}[||x_{T}-\hat{x}_{T}^{(h)}||^{2}_{2}]&=E_{p(x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})}[\text{Var}_{p(x_{T}|x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})}(x_{T})],\\ \hat{x}_{T}^{(h)*}(x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})&=E_{p(x_{T}|x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})}[x_{T}],\end{split}$ (19)
@@
$\begin{split}\min_{\hat{x}_{T}^{(h)}}E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})}[||x_{T}-\hat{x}_{T}^{(h)}||^{2}_{2}]&=E_{p(x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})}[\text{Var}_{p(x_{T}|x_{s_{1}},\cdot\cdot\cdot,x_{s_{l}})}(x_{T})]\\ &\leq E_{p(x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[\text{Var}_{p(x_{T}|x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}(x_{T})]=\min_{\hat{x}_{T}^{(g)}}E_{p(x_{T},x_{s_{1}},\cdot\cdot\cdot,x_{s_{k}})}[||x_{T}-\hat{x}_{T}^{(g)}||^{2}_{2}].\end{split}$ (20)
@@ Appendix C Hyperparameters
Appendix D PSNR for Reconstruction
Table 10: PSNR of reconstruction for the last frame of pseudo videos produced by C-ViViT trained on pseudo videos constructed from CIFAR10 and CelebA images. 1-frame results are obtained from their image counterparts VQ-VAE trained on original CIFAR10 and CelebA images.
CIFAR10 CelebA
1-frame 8-frame 18-frame 1-frame 8-frame 18-frame
Reconstruction 17.20 21.86 25.89 20.52 24.66 29.68

### v3

@@ 2.3 Improving Image Generation via Pseudo Video Generation
With the concept of pseudo videos and their effectiveness in diffusion models, we are interested in the following open generic question in this paper: Is it possible to improve other types of image generative models by jointly modelling the distribution of the original image and its corresponding pseudo video which contains self-supervised information? The answer is affirmative. In this paper, we show empirical evidence of the advantages of pseudo videos on two types of generation models, namely improving VQVAE (Van Den Oord et al., 2017) (Section 3) and DDPM (Nichol & Dhariwal, 2021) (Section 4.2) with Phenaki (Villegas et al., 2022) and Video Diffusion (Harvey et al., 2022) trained on pseudo videos, respectively. Moreover, we provide theoretical arguments favouring the use of more expressive ways of creating pseudo videos in the autoregressive video generation framework (Section 4.1), beyond the first-order Markov strategy as typically used in the forward process of standard diffusion models.
@@ 3 Improved Reconstruction and Generation in VQ-VAE with Pseudo Videos
Datasets. We create 8-frame and 18-frame pseudo videos using images from two benchmark datasets, CIFAR10 (32 $\times$} 32) (Krizhevsky et al., 2009) and CelebA (64 $\times$} 64) (Liu et al., 2015), with the blurring technique from Bansal et al. (2023). To create 8-frame pseudo videos, we blur the images recursively 7 times with a Gaussian kernel of size $\times11$ and standard deviation growing exponentially at the rate of 0.05. For 18-frame pseudo videos, we blur the images 17 times with same Gaussian kernel but with a standard deviation growing exponentially at the rate of 0.01. The pseudo videos are organized such that the last frames are the original target images and the first frames are the blurriest images. Figure 3 shows an example of such a pseudo video. We use 1-frame to denote images generated by the original image autoencoder/generative model trained on the original target images.
@@
Evaluation Metric. We compute Frechet Inception Distance (FID) (Heusel et al., 2017) with 50k samples to evaluate the quality of images, either from the last-frames of videos generated by video models trained on pseudo videos, or images generated by image models trained on the original target images. We also report Peak Signal-to-noise Ratio (PSNR $\uparrow$) as an additional metric for reconstruction result.
@@
Results. Table 2 shows the last-frame FID and PSNR of C-ViViT for reconstruction and that of VideoGPT and Phenaki for AR and masked generation on CIFAR10 and CelebA, respectively. The 1-frame results correspond to the performance of their image model counterparts (i.e., VQVAE for image reconstruction, and ImageGPT and MaskGit for AR and masked image generation, respectively). We observe that pseudo videos indeed help improve the training of the C-ViViT as the reconstruction quality of the last frame is significantly improved with a few more frames. We show reconstructed CIFAR10 and CelebA images from different C-ViViT models in Figures LABEL:fig:recon_cifar10 and LABEL:fig:recon_celeba, respectively. It can be seen that for both datasets, the reconstructed images trained with 1-frame models are relatively smoothed and this issue is resolved by using pseudo videos which produce sharper images. Quantitatively, reconstruction FID improves as more frames are used. Pseudo videos also improve the image generation performance compared to the 1-frame results. Interestingly, we see a diminishing return as we include more frames. For CelebA images, 18-frame pseudo videos help achieve the best image generation performance for both AR and latent masked generation, but 8-frame models achieve a comparable performance. For CIFAR10 images, 8-frame pseudo videos result in better image generation performance than 18-frame pseudo videos, which suggests the latent tokens have a more complex prior distribution in order to reconstruct model pseudo videos with 18 frames well and therefore this prior is more difficult for VideoGPT or Phenaki to capture. In summary. while pseudo videos help improve the performance, the optimal number of frames may depend on the dataset and the augmentation strategy. This diminishing return is not a severe issue in practice since practitioners may prefer to improve the generation with just a few more pseudo frames to avoid introducing a high computational cost.
@@
Table 2: Last-frame FID/PSNR of images produced by C-ViViT (reconstruction), VideoGPT (AR generation) and Phenaki (latent masked generation) trained on pseudo videos constructed from CIFAR10 and CelebA images. 1-frame results are obtained from their image counterparts VQVAE (reconstruction), ImageGPT (AR generation) and MaskGit (latent masked generation) trained on original CIFAR10 and CelebA images.
CIFAR10 CelebA 1-frame 8-frame 18-frame 1-frame 8-frame 18-frame Reconstruction (FID) 24.53 13.81 11.26 24.62 5.72 2.27 Reconstruction (PSNR) 19.97 21.86 25.89 20.52 24.66 29.68 AR Generation (FID) 72.06 54.60 69.23 32.98 30.19 28.08 Latent Masked Generation (FID) 49.27 35.50 47.65 27.34 16.87 16.66
@@ 4.1 Is First-order Markov Chain the Optimal Choice for Creating Pseudo Videos?
Moreover, the non-optimality of creating pseudo video via first-order Markov chain becomes clear: the first-order Markov data augmentation implies that $p(x_{T}|x_{T-1})=p(x_{T}|x_{T-1},x_{T-2})$ and consequently $\mathcal{L}_{2}^{*}=\mathcal{L}_{1}^{*}$. More specifically, for strict inequality in Eq 14, we need to avoid $p(x_{T}|x_{T-1})=p(x_{T}|x_{T-1},x_{T-2})$, which is equivalent to avoiding the use of either first-order Markov chain $x_{T}\rightarrow x_{T-1}\rightarrow x_{T-2}$ or $x_{T}\leftarrow x_{T-1}\rightarrow x_{T-2}$. This analysis is informative for us to design better pseudo videos, for example through data augmentation with higher-order Markov chains. We formalize the above informal reasoning into Theorem 4.1 and provide the formal proof in Appendix A.
@@
[…]
@@ 4.2 Experiments
Datasets. We create 4-frame and 8-frame pseudo videos using images from CIFAR10 (32 $\times$} 32) and CelebA (64 $\times$} 64). We use Gaussian noise as data augmentation and we consider two strategies:
@@ 6 Conclusion and Discussions
Reproducibility Statement
Code for experiments in Section 3 can be found at https://github.com/Wenlin-Chen/pseudo-video-gen, which is adapted from https://github.com/lucidrains/phenaki-pytorch. The experiments in Section 4.2 are carried out directly using the official implementation of Flexible Video Diffusion Model (FDM) (Harvey et al., 2022) and Improved DDPM (Nichol & Dhariwal, 2021), from https://github.com/plai-group/flexible-video-diffusion-modeling and https://github.com/openai/improved-diffusion.
Appendix A Proof of Theorem 4.1
@@
[…]
@@
[…]
@@
[…]
@@
[…]